OpenStudy (anonymous):

Find the unit vector in the direction opposite to v=(-5,1).

jimthompson5910 (jim_thompson5910):

First find the vector w that is pointing in the opposite direction of v

jimthompson5910 (jim_thompson5910):

Then find the unit vector y such that y is pointing in the same direction as w and |y| = 1

OpenStudy (anonymous):

so would w=(5,-1 and |y|= sort(26)

jimthompson5910 (jim_thompson5910):

w = (5,-1), good

jimthompson5910 (jim_thompson5910):

the length of w is |w| = sqrt(26)

jimthompson5910 (jim_thompson5910):

you divide each piece of vector w by its length to get the unit vector y (which points in the same direction as w)

OpenStudy (anonymous):

so the opposite direction unit vector for v is 5/sqrt(26)+(-1)/sqrt(26)

jimthompson5910 (jim_thompson5910):

You can write it as \[\Large \vec{y} = \frac{5}{\sqrt{26}}i+\frac{-1}{\sqrt{26}}j\] or as \[\Large \vec{y} = \left<\frac{5}{\sqrt{26}}, \frac{-1}{\sqrt{26}}\right>\] the angle brackets signify we have a vector (and not just a point or ordered pair)

jimthompson5910 (jim_thompson5910):

Optionally you can rationalize the denominators, but I've seen many cases where the book doesn't rationalize the denominator

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome